OpenStudy (anonymous):

find the indicated outputs for f(x)=4x to the second power-5x, f(0)= f(-1)= f(2)=

OpenStudy (anonymous):

I don't understand the question

OpenStudy (anonymous):

f(x) = (4x)^(-5x) ?

jimthompson5910 (jim_thompson5910):

f(x)=4x^2-5x f(0)=4(0)^2-5(0) f(0)=4(0)-5(0) f(0)=0-0 f(0) = 0 ---------------------------- f(x)=4x^2-5x f(-1)=4(-1)^2-5(-1) f(-1)=4(1)-5(-1) f(-1)=4+5 f(-1)=9 ---------------------------- f(x)=4x^2-5x f(2)=4(2)^2-5(2) f(2)=4(4)-5(2) f(2)=16-10 f(2)=6

OpenStudy (anonymous):

find the indicated outputs for f(x)=4x second power - 5x. f(0)= f(-1)= f(2)=